# EAMA : Entity-Aware Multimodal Alignment Based Approach for News Image Captioning

### v4

@@ 3.3. Self-Supplemented Generation
To enhance news image caption generation, models rely on concise and pertinent textual information. Previous research often employs a specific length limitation from the start of news articles or paragraphs surrounding the provided image due to token limitations. In our method, the aligned MLLM initially extracts sentences which contain entities relevant to the given news image from the entire news article. Because the aligned MLLM is trained on the Entity-Aware Sentence Selection task during the alignment training stage, this extraction stage doesn’t incur additional model or training costs.
@@
Once we acquire entity-related sentences, we seamlessly integrate them with the default textual input context. To ensure coherence, we eliminate duplicated sentences and amalgamate the rest sentences into the final textual input context, maintaining their original order in the news article. To prevent the final textual input context from becoming overly long, we impose limits on the number of added supplemental sentences and restrict the overall length of the final input context.
Method Train BLEU-4 ROUGE CIDEr Named entities
Setting P R
Base OSFT 10.05 25.45 75.95 27.39 30.37
Base Fully_SFT 10.88 26.60 82.70 29.22 31.32
Align($\alpha\mathcal{L}^{\mathrm{SENT}}+\gamma\mathcal{L}^{\mathrm{CAP}}$) Fully_SFT 10.92 26.68 83.60 29.42 31.45
Align($\beta\mathcal{L}^{\mathrm{ENT}}+\gamma\mathcal{L}^{\mathrm{CAP}}$) Fully_SFT 10.97 26.87 84.28 29.26 31.58
Align($\mathcal{L}^{\mathrm{SENT}}+\mathcal{L}^{\mathrm{ENT}}+\mathcal{L}^{\mathrm{CAP}}$) Fully_SFT 10.58 26.18 81.94 29.61 30.69
Align($\alpha\mathcal{L}^{\mathrm{SENT}}+\beta\mathcal{L}^{\mathrm{ENT}}+\gamma\mathcal{L}^{\mathrm{CAP}}$) Fully_SFT 10.80 26.80 84.45 29.15 31.45
EAMA Fully_SFT 10.94 27.00 85.61 29.32 31.85
@@
Table 3. Ablation study results on the NYTimes800k dataset. Base indicates finetuning InstructBLIP with the news image caption task. Align refers to aligning InstructBlip with different alignment tasks. $\alpha$, $\beta$ and $\gamma$ represent the loss weights in Equation 7. EAMA denotes our entire approach. We omit $\alpha$, $\beta$ and $\gamma$ when they are all 1. In Train Setting, ”OSFT” denotes finetuning the V-L Connector while keeping the LLM frozen, Fully_sft denotes fully finetuning both V-L Connector and LLM in InstructBLIP.
4. Experiment Setup
4.1. Datasets & Metrics
@@
In our experiments, we utilize two publicly available large-scale News Image Captioning datasets: GoodNews (Biten et al., 2019), comprising 421K, 18K, and 23K samples for train, validation, and test splits respectively, and NYTimes800k (Tran et al., 2020), containing 763K, 8K, and 22K samples for train, validation, and test splits, sourced from The New York Times API.
@@
We use BLEU-4 (Papineni et al., 2002), ROUGE (Lin, 2004) and CIDEr (Vedantam et al., 2015) as automatic evaluation metrics. Following previous research (Tran et al., 2020), we calculate these metrics with the COCO caption evaluation toolkit [^1] https://github.com/tylin/coco-caption. According to previous researches (Tran et al., 2020; Zhou et al., 2022a; Zhang and Wan, 2023), we prioritize CIDEr as the most crucial metric because it places greater emphasis on uncommon words, which often include specific named entities. Additionally, we evaluate the Precision and Recall of named entities. To accomplish this, we employ SpaCy to recognize entities in both generated captions and gold captions, counting the number of exact matches, which necessitates identical text and entity labels.
4.2. Implementation Details
@@
When finetuning MiniGPT-v2, InstructBLIP, and LLaVA-v1.5, adhering to the official finetune scripts, we fine-tune the linear projection layer of MiniGPT-v2, the Q-Former, and Fully Connected Layer of InstructBLIP, and fully fine-tune the MLP layer and LLM of LLaVA-v1.5 on checkpoints of 7B. Our training and inference prompts can be found in Appendix A, which remain unchanged among our experiments.
When aligning our model with the two designed Entity-Aware alignment tasks, we set the values of $\alpha$, $\beta$, and $\gamma$ in Equation (7) as follows: 0.5, 0.25, and 1 respectively for the GoodNews dataset, and 0.25, 0.75, and 1 respectively for the NYTimes800k dataset. These values are determined based on their performance on the validation dataset. For more implementation details, please refer to Appendix B.
Method Train Textual Input of NIC Textual Input of NIC BLEU-4 ROUGE CIDEr Named entities
Setting Training Inference P R
Base Fully_SFT Origin Origin 10.88 26.60 82.70 29.22 31.32
Base Fully_SFT Full Full 10.76 26.18 83.17 27.71 30.03
Base Fully_SFT Origin Origin(Longer) 10.60 26.28 81.75 28.95 30.57
Base Fully_SFT Origin Oracle 11.24 27.70 87.56 33.42 34.80
Align Fully_SFT Origin Origin 10.80 26.80 84.45 29.15 31.45
Align Fully_SFT Origin Origin(Longer) 10.75 26.70 83.50 29.25 31.16
EAMA Fully_SFT Origin Supplemented 10.94 27.00 85.61 29.32 31.85
Table 4. Analysis of Textual Input with InstructBLIP in News Image Captioning (NIC) on NYTimes800k dataset. Base indicates finetuning InstructBLIP with the NIC task. Align means align InstructBLIP with all three alignment tasks through Balance Training Strategy. Fully_sft denotes finetuning both the V-L Connector and LLM. Origin denotes the traditional textual input of NIC. Full denotes the entire news article. Supplemented denotes fusing the supplementary sentences with Origin with a limited length. Origin(Longer) denotes obtaining textual input of the same maximum length as Supplemented but using traditional setting. Oracle denotes applying the sentences containing the same entities as captions, serving as an ideal upper bound.
5. Experimental Results
5.1. Performance of MLLMs on NIC
@@
LLaVA-v1.5(Liu et al., 2023a), MiniGPT-v2(Chen et al., 2023) and InstructBLIP(Dai et al., 2023) are three common MLLMs. We first test them under zero-shot setting and present results in Table 1. The results presented in Table 1 further confirm our initial hypothesis that MLLMs trained on general multimodal data may struggle to generate informative news image captions effectively. Notably, the Recall of named entities is consistently lower than the Precision of named entities on MLLMs, especially on LLaVA-v1.5, which means MLLMs tend to not generate entities under zero-shot setting. In fact, the average number of generated entities (whether correct or incorrect) by LLaVA-v1.5 are 1.22 and 1.24 on GoodNews and NYTimes800k datasets, while the average number of entities in references are 3.43 and 3.19 on these two datasets respectively. This substantial disparity lends further support to our observations.
@@
Subsequently, we conduct straightforward finetuning of these three MLLMs with their official scripts on news image captioning dataset. Following their official scripts, we finetune the V-L Connector and LLM component of LLaVA-v1.5 and the V-L Connector component of MiniGPT-v2 and InstructBLIP. We refer to this training setting as OSFT. According to the results in Table 1, this simple finetuning approach yields promising performance improvements on both the GoodNews and NYTimes800k datasets. Thus we posit that MLLMs demonstrate competence in multimodal Entity-Aware tasks when trained appropriately on multimodal data rich in entity information. Besides, we observe that the OSFT InstructBLIP obtains better performance than LLaVA-v1.5 and MiniGPT-v2 on both dataset, especially on NYTimes800k which contains longer news articles. We attribute this to the V-L Connector module of InstructBLIP, which compresses visual information into 32 tokens. This enables InstructBLIP to accommodate more textual input compared to other MLLMs, making InstructBLIP the most suitable MLLM among all three MLLMs for news image captioning. Therefore, without loss of generality, we select InstructBLIP for further investigation and experimentation. More analysis on entity generation is provided in Appendix C.
5.2. Comparison Results of Different Methods
@@
Table 2 presents a comparison of the performance of our proposed approach (EAMA) and the baseline models, including the OSFT InstructBLIP settings on two news image captioning datasets. The results in Table 2 show that EAMA achieves the best performance across all metrics. Compared to the previous state-of-the-art (SOTA) methods (Rajakumar Kalarani et al. (2023) for GoodNews and Rajakumar Kalarani et al. (2023) for NYTimes800k), our model gains 16.06 and 14.78 points in CIDEr score on the GoodNews and NYTimes800k datasets, respectively. Furthermore, our method outperforms the OSFT InstructBLIP approach by 10.36 and 9.66 points in CIDEr score on the GoodNews and NYTimes800k datasets, respectively. In addition to the superior performance on automatic evaluation metrics, our method also achieves the best results in entity generation, obtaining high precision and recall simultaneously.
5.3. Ablation Study
@@
We conducted ablation experiments on the NYTimes800k dataset to study the impact of different components of our approach. In addition to the standard fine-tuning of InstructBLIP, we also fully fine-tuned both the Vision-Language Connector and the Large Language Model, which we refer to as (Fully_SFT). The results in Table 3 show that fully fine-tuning the Large Language Model is crucial for InstructBLIP to generate high-quality news image captions. We hypothesize that handling entity information, such as people’s names, using multimodal information is a very challenging task, distinct from other common multimodal data, making fine-tuning the Vision-Language Connector alone insufficient.
@@
We also conducted ablation studies to demonstrate the effectiveness of the Entity-Aware Sentence Selection and Entity Selection tasks individually. In Table 3, Align ($\alpha\mathcal{L}^{\mathrm{SENT}}+\beta\mathcal{L}^{\mathrm{ENT}}+\gamma\mathcal{L}^{\mathrm{CAP}}$) means aligning the model on the Entity-Aware Sentence Selection, Entity Selection, and News Image Captioning tasks simultaneously, with the corresponding loss weights $\alpha$, $\beta$, and $\gamma$ as defined in Equation 7 of the Balance Training Strategy. Align ($\alpha\mathcal{L}^{\mathrm{SENT}}+\gamma\mathcal{L}^{\mathrm{CAP}}$) and Align ($\beta\mathcal{L}^{\mathrm{ENT}}+\gamma\mathcal{L}^{\mathrm{CAP}}$) follow the same pattern. From the ablation study results, we observe that both alignment tasks are effective when used separately. The direct impact of the Entity Selection task is more pronounced, but applying Sentence Selection can further improve the performance. Hence, both alignment tasks are effective and necessary components of our alignment framework.
We also investigated the impact of the Balance Training Strategy. Align ($\mathcal{L}^{\mathrm{SENT}}+\mathcal{L}^{\mathrm{ENT}}+\mathcal{L}^{\mathrm{CAP}}$) aligns the model without the Balance Training Strategy, which means $\alpha$, $\beta$, $\gamma$ are all set to 1. The results indicate that the Balance Training Strategy can better align the InstructBLIP model with the alignment tasks and achieve improved performance on the CIDEr metric and Recall of entities.
@@
EAMA refers to our approach, which first aligns the InstructBLIP model with all three tasks using the Balance Training Strategy, and then applies Self-Supplemented Generation during the inference stage when generating news image captions. With the addition of Self-Supplemented Generation, EAMA further improves the performance from a CIDEr score of 84.45 to 85.61, and also achieves better generation of entities.
Figure 3. An example of news image caption generation. We present the news image, news article and the news image caption of one sample from NYTimes800k dataset. The news image captions are respectively, generated in: (1) InstructBLIP in zero-shot setting. (2) LLaVA-v1.5 in zero-shot setting. (3) GPT-4V in zero-shot setting. (4) InstructBLIP (OSFT) on NYTimes800k. (5) LLaVA-v1.5 (OSFT) on NYTimes800k. (6) Our method, EAMA. We highlight Entities occurred in reference in red.
5.4. Analysis on Textual Input
@@
In our experiments on the textual input for the news image captioning task, we observe that most previous methods limit the length of textual context due to the maximum token constraint. As indicated in (Tran et al., 2020), the average length of news articles in the NYTimes800k dataset is 974 tokens. Given that InstructBLIP has a maximum of 2048 tokens, with only 32 of them being visual tokens, we attempt to use entire news articles as input during both training and inference stages. This approach aims to leverage more textual information into the MLLM, potentially enhancing performance. However, our experiments reveal that the performance only increases marginally, from 82.70 to 83.17 in CIDEr score. This modest improvement is not satisfactory, considering that using such long textual input nearly doubles the time and memory costs.
@@
To delve deeper, we conduct an additional experiment by testing the model with oracle-selected sentences as textual input. These oracle sentences averagely contain only 14% of the words from entire news articles. They are chosen from news articles and contain the same entities as reference captions. While this scenario is ideal and impossible in practice since we cannot access reference captions during inference, the results indicate that the oracle textual input can achieve competitive performance. Therefore, compared to simply expanding the length of textual input, we believe that selecting entity-related sentences for supplementation is more effective.
@@
5.5. Case Study
We randomly select a sample from NYTimes800k dataset and generate captions using different models, as illustrated in Figure 3. From the generated captions, it’s evident that MLLMs under zero-shot settings can only produce common entity information such as ”The United States” or ”England”, even when provided with the news image and associated text. Even the most powerful MLLM, GPT-4V (OpenAI et al., 2023) fails to generate specific entities related to the news image and mentioned in the news article, such as ”Julie Ertz” and ”Lucy Bronze.” After OSFT on NYTimes800k dataset, InstructBLIP still cannot generate the names of people in the news image. LLaVA-v1.5 successfully generates ”Bronze” but overlooks ”England.” In contrast, our method generates both common entities and the names of people mentioned in the news article, further affirming the effectiveness of our approach in entity recognition. Details about our case example can be found in Appendix D.
6. Conclusion
In this work, we reveal that MLLMs trained on conventional multimodal data struggle with generating entities in news image captioning tasks. Simply fine-tuning MLLMs proves insufficient for enabling them to effectively handle multimodal entity information. To tackle this challenge, we introduce EAMA, which aligns the MLLM with two entity-aware alignment tasks and employs a balanced training strategy. Subsequently, we utilize the aligned MLLM to supplement the textual input context, as we observe that generating high-quality news image captions necessitates entity-related and concise input texts. Our proposed method, EAMA, not only achieves superior performance on automatic evaluation metrics but also demonstrates the best ability to generate entities in news image captioning tasks.

### v5

@@ 3.3 Self-Supplemented Generation
To enhance news image caption generation, models rely on concise and pertinent textual information. Previous research often employs a specific length limitation from the start of news articles or paragraphs surrounding the provided image due to token limitations. In our method, the aligned MLLM initially extracts related entities and sentences which contain entities relevant to the given news image from the entire news article. Because the aligned MLLM is trained on the Entity-Aware Sentence Selection task and Entity Selection task during the alignment training stage, this extraction stage doesn’t incur additional modules.
@@
Once we acquire related entities and entity-related sentences, we seamlessly integrate them with the default textual input context. To ensure coherence, we eliminate duplicated sentences and combine the rest sentences into the final textual input context, maintaining their original order in the news article. To prevent the final textual input context from becoming overly long, we restrict the overall length of the final input context. More details can be seen in Appendix B.
Method BLEU-4 METEOR ROUGE CIDEr Named entities
P R
GoodNews Avg+CtxIns(Biten et al., 2019) 0.89 4.37 12.20 13.10 8.23 6.06
ICECAP(Hu et al., 2020) 1.96 6.01 15.70 26.08 / /
Tell(Tran et al., 2020) 6.05 10.30 21.40 53.80 22.20 18.70
JoGANIC(Yang et al., 2021) 6.83 11.25 23.05 61.22 26.87 22.05
NewsMEPZhang et al. (2022) 8.30 12.23 23.17 63.99 23.43 23.24
Rule-driven(Xu et al., 2024b) 8.18 12.50 23.56 71.58 25.51 23.68
Rajakumar Kalarani et al. (2023) 7.14 11.21 24.30 72.33 24.37 20.09
Qu et al. (2024) 8.60 12.39 23.38 71.96 24.30 25.54
Xu et al. (2024a) 8.49 12.88 26.22 83.52 30.19 26.57
InstructBLIP (OSFT) 9.53 13.54 25.61 78.03 25.89 27.33
EAMA 10.04 13.95 27.06 87.70 27.58 28.92
NYTimes800k Tell(Tran et al., 2020) 6.30 10.30 21.70 54.40 24.60 22.20
JoGANIC(Yang et al., 2021) 6.79 10.93 22.80 59.42 28.63 24.49
NewsMEP(Zhang et al., 2022) 9.57 13.02 23.62 65.85 26.61 28.57
Rule-driven(Xu et al., 2024b) 9.41 13.10 24.42 72.29 28.15 28.80
Rajakumar Kalarani et al. (2023) 7.54 11.27 23.28 66.41 28.11 23.25
Qu et al. (2024) 9.24 12.57 23.44 71.65 26.88 28.59
Xu et al. (2024a) 9.07 13.17 26.48 83.72 32.38 30.08
InstructBLIP (OSFT) 10.05 13.63 25.45 75.95 27.39 30.37
EAMA 11.03 14.22 27.15 87.00 29.79 32.24
Table 1: Experimental results on GoodNews and NYTimes800k datasets compared with other models. P and R denote Precision and Recall of generated Named entities. InstructBLIP(OSFT) refers to InstructBLIP trained with official scripts on 7B checkpoints. We directly cite the results of models which are not MLLMs from researches by Rajakumar Kalarani et al. (2023); Qu et al. (2024).
Method BLEU-4 METEOR ROUGE CIDEr Named entities
P R
InstructBLIP 10.05 13.63 25.45 75.95 27.39 30.37
Base 10.88 14.09 26.60 82.70 29.22 31.32
Align(SENT $+$ CAP) 10.92 14.08 26.68 83.60 29.42 31.45
Align(ENT $+$ CAP) 10.97 14.20 26.87 84.28 29.26 31.58
Align(SENT $+$ ENT $+$ CAP) 10.80 14.08 26.80 84.45 29.15 31.45
EAMA 11.03 14.22 27.15 87.00 29.79 32.24
@@
Table 2: Ablation study results on the NYTimes800k dataset. InstructBLIP refers to OSFT InstructBLIP. Base indicates fully finetuning InstructBLIP on the news image caption task. Align refers to aligning InstructBLIP with different alignment tasks. EAMA denotes our entire approach. According to Sec 4.1, CIDEr is the most crucial metric because it places greater emphasis on uncommon words, which often include specific named entities.
4 Experiment Setup
4.1 Datasets & Metrics
@@
In our experiments, we utilize two publicly available large-scale News Image Captioning datasets [^2] We don’t have the official test split of VisualNews dataset, so we conduct experiments on the same two datasets with Qu et al. (2024) and Xu et al. (2024a).: GoodNews (Biten et al., 2019), comprising 421K, 18K, and 23K samples for train, validation, and test splits respectively, and NYTimes800k (Tran et al., 2020), containing 763K, 8K, and 22K samples for train, validation, and test splits.
@@
We use BLEU-4 (Papineni et al., 2002), METEOR (Denkowski and Lavie, 2014), ROUGE (Lin, 2004) and CIDEr (Vedantam et al., 2015) as automatic evaluation metrics, following previous research (Tran et al., 2020)  [^3] These metrics are calculated with the COCO caption evaluation toolkit https://github.com/tylin/coco-caption. According to previous researches (Tran et al., 2020; Zhou et al., 2022a; Zhang and Wan, 2023), CIDEr is the most crucial metric because it places greater emphasis on uncommon words, which often include specific named entities. Additionally, we evaluate the Precision and Recall of named entities. To accomplish this, we employ SpaCy to recognize entities in both generated captions and gold captions and count the number of exact matches.
4.2 Implementation Details
@@
When simply finetuning MiniGPT-v2, InstructBLIP, and LLaVA-v1.5, adhering to the official scripts, we fine-tune the linear projection layer of MiniGPT-v2, the Q-Former, and Fully Connected Layer of InstructBLIP, and both the MLP layer and LLM of LLaVA-v1.5 on checkpoints of 7B. Our training and inference prompts can be found in Appendix A, which remain unchanged among our experiments. All training hyper-parameters follow their default configurations: the initial learning rates for InstructBLIP and LLaVA-v1.5 are set to $2\times 10^{-5}$, while the initial learning rate for MiniGPT-v2 is set to $1\times 10^{-5}$. During inference, we use a beam size of 5 in MLLMs. Mixed precision training with bp16 is employed to accelerate the training process. Our model takes approximately 2-3 days to train on four Nvidia A40 GPUs. More implementation details can be found in Appendix B.
5 Experimental Results
5.1 Performance of MLLMs on NIC
LLaVA-v1.5(Liu et al., 2023a), MiniGPT-v2(Chen et al., 2023) and InstructBLIP(Dai et al., 2023) are three common MLLMs. We show the representative results in Figure 3 and full evaluation results are shown in Table 3 in Appendix C.
@@ 5.1 Performance of MLLMs on NIC
Firstly, the results under zero-shot setting confirm our initial hypothesis that MLLMs trained on general multimodal data struggle to generate informative news image captions. Notably, the Recall of named entities is lower than the Precision of named entities on MLLMs, especially for LLaVA-v1.5, which means MLLMs tend not to generate entities under zero-shot setting. In fact, the average number of generated entities (whether correct or not) by LLaVA-v1.5 are 1.22 and 1.24 on GoodNews and NYTimes800k datasets, while the average number of entities in references are 3.43 and 3.19 on these two datasets respectively. This substantial disparity lends further support to our observations.
@@
Subsequently, we conduct straightforward finetuning of these three MLLMs with their official scripts on news image captioning dataset. Following their official scripts, we finetune the V-L Connector and LLM component of LLaVA-v1.5 and the V-L Connector component of MiniGPT-v2 and InstructBLIP. We refer to this training setting as OSFT. According to the results in Figure 3, this simple finetuning approach yields promising performance improvements on both the GoodNews and NYTimes800k datasets. Thus we believe that MLLMs are capable of multimodal Entity-Aware tasks when trained appropriately on multimodal data rich in entity information. Besides, we observe that the OSFT InstructBLIP obtains better performance than LLaVA-v1.5 and MiniGPT-v2 on both dataset, especially on NYTimes800k which contains longer news articles. We attribute this to the V-L Connector module of InstructBLIP, which compresses visual information into 32 tokens. This enables InstructBLIP to accommodate more textual input compared to other MLLMs, making InstructBLIP the most suitable MLLM among all three MLLMs for news image captioning. Therefore, we select InstructBLIP for further investigation and experimentation.
5.2 Comparison Results of Different Methods
@@
Table 1 presents a comparison of the performance of our proposed approach (EAMA) and the baseline models, including the OSFT InstructBLIP settings on two news image captioning datasets. The results in Table 1 show that our method outperforms the OSFT InstructBLIP approach by 9.67 and 11.05 points in CIDEr score on the GoodNews and NYTimes800k datasets, respectively. Furthermore, our method EAMA achieves the best performance across all automatic metrics. Compared to the previous state-of-the-art (SOTA) method (Xu et al. (2024a) for GoodNews and NYTimes800k), our model gains 4.18 and 3.28 points in CIDEr score on the GoodNews and NYTimes800k datasets, respectively. In addition to the superior performance on automatic evaluation metrics, our method also achieves the highest recall of named entities simultaneously. We also provide analysis on entity generation in Appendix E and a case study in Appendix F to offer a comprehensive understanding of EAMA.
5.3 Ablation Study
@@
We conduct ablation experiments on the NYTimes800k dataset to study the impact of different components of our approach. In addition to the OSFT of InstructBLIP, we also fully fine-tuned both the Vision-Language Connector and the Large Language Model, which we refer to as Base. The results in Table 2 show that fully fine-tuning the Large Language Model is crucial for InstructBLIP to generate high-quality news image captions. We hypothesize that handling entity information using multimodal information is very challenging for its distinction from other common multimodal data, making fine-tuning the Vision-Language Connector alone insufficient.
@@
We also conducted ablation studies to demonstrate the effectiveness of the Entity-Aware Sentence Selection and Entity Selection tasks individually. In Table 2, Align(SENT $+$ ENT $+$ CAP) means aligning the model on the Entity-Aware Sentence Selection, Entity Selection, and News Image Captioning tasks simultaneously. Align(SENT $+$ CAP) and Align(ENT $+$ CAP) follow the same pattern. From the ablation study results, we observe that both alignment tasks are effective when used separately. Both alignment tasks are effective and necessary components of our alignment framework.
@@
EAMA refers to our approach, which first aligns the InstructBLIP model with all three tasks, and then applies Self-Supplemented Generation during the inference stage when generating news image captions. With the addition of Self-Supplemented Generation, EAMA further improves the performance from a CIDEr score of 84.45 to 87.00, and also achieves better generation of entities.
5.4 Analysis on Textual Input
@@
We would also like to investigate the influence of different textual input on news image captioning task. The full results are shown in Table 4 in Appendix D. In our experiments on the textual input for the news image captioning task, we observe that most previous methods limit the length of textual context due to the maximum token constraint. As indicated in (Tran et al., 2020), the average length of news articles in the NYTimes800k dataset is 974 words. Given that InstructBLIP has a maximum of 2048 tokens, with only 32 of them being visual tokens, we attempt to use entire news articles as input during both training and inference stages. This approach aims to leverage more textual information into the MLLM, potentially enhancing performance. However, our experiments reveal that the performance only increases marginally, from 82.70 to 83.17 in CIDEr score. This modest improvement is not satisfactory, considering that using such long textual input nearly doubles the time and memory costs during training and inference stages.
@@
To research deeper, we conduct an additional experiment by testing the model with oracle-selected sentences as textual input. These oracle sentences averagely contain only 14% of the words from entire news articles. They are chosen from news articles and contain the same entities as reference captions. We also test the model with oracle-selected entities which occur both in the news article and the caption. To obtain the oracle upper bound, We test the model with input textual context supplemented with both oracle-selected sentences and entities. While these scenarios are ideal and impossible in practice since we cannot access reference captions during inference, the results indicate that the oracle supplemented textual input can achieve competitive performance. Therefore, compared to simply expanding the length of textual input, we believe that selecting related entities and entity-related sentences for supplementation is more effective.
@@
6 Conclusion
In this work, we reveal that MLLMs trained on conventional multimodal data struggle with generating entities in news image captioning task. Simply fine-tuning MLLMs proves insufficient for enabling them to effectively handle multimodal entity information. To tackle this challenge, we introduce EAMA, which aligns the MLLM with two entity-aware alignment tasks. Subsequently, we utilize the aligned MLLM to supplement the textual input context, as we observe that generating high-quality news image captions necessitates entity-related and concise textual input context. Our proposed method, EAMA, not only achieves superior performance on automatic evaluation metrics but also obtains the competitive result of entity generation in news image captioning task.
7 Limitations
We conduct our experiments mainly on three common MLLMs. The conclusion will be more general if we test more MLLMs. Our method requires big GPU memory both on training and testing stages, and is slower than previous models. With the Optimization technique such as flash-attention provided, the speed of our method will become faster. The data of our alignment tasks are extracted only from news image captioning tasks. The knowledge of news is updating rapidly. Our method still need to align the model on data which contains new updated knowledge to obtain the best performance.
8 Ethics Statement
The generation of our method is highly related to the pretrained multimodal checkpoints and the alignment task data. They may both introduce bias and hallucination into the generated captions which may cause inaccuracy news report. We recommend to manual check the results before actually using in the real situation.